R1 = 1kΩ
R2 = 3kΩ
R3 = R4 = 1kΩ
Voffset = 1V
Vin = 2V

Determine the value of the voltage in Vout.

R1 R2 ww o Vout Vin o +15V opamp U1 15V R4 R3 Voffset
Answer #1

IKA RI ww Зкл VAR2 Vin Vout ASIT )К^ Iкл Vc V V- R3 + V 0.SV V V= 0-SV V2 V1 VA-Vout - O VA V = 0 .SV 0. S-Vout O S- 2 20 3 K (M01XE991) 3.333Xio7 vout (-sx Vout 4V